A group of graduate students completed a cohort study that followed 785 drivers for a period of one year to explore the relationship between texting while driving and motor vehicle accidents. Three hundred thirty (330) of the drivers texted while driving in the past year. Four hundred fifty-five drivers did not text while driving during the same time. Thirty (30) motor vehicle accidents occurred in the texting group, and 5 motor vehicle accidents occurred in the non-texting group. Set up a 2 x 2 contingency table reflecting this data.
Answer #1
Accident occured No accident Total
Drivers exposed to texting 30(a) 300(b) 330 Exposure rate among texting group(a/a+c)x100=30/35 x100=85%
Drivers not texting 5 (c) 450(d) 455 Exposure rate among non texting group=(b/b+d)x100=300/750x100=40%
Total 35 750 785 chi square =(axd--bxc)2 x (into) total no of drivers (785)/(divided by)(330x455x35x750) = .........P value